write 3 different series each with 5 terms and each with a sum of 100.
Found 3 solutions by Edwin McCravy, AnlytcPhil, Theo:
Answer by Edwin McCravy(20056)
Answer by AnlytcPhil(1806)


-10+5+20+35+50 = 100, is an arithmetic series with 1st term -10
 and common difference 15.

-8+6+20+34+48 = 100, is an arithmetic series with 1st term -8
 and common difference 14.

-6+7+20+33+46 = 100, is an arithmetic series with 1st term -6
 and common difference 13.

-4+8+20+32+44 = 100, is an arithmetic series with 1st term -4
 and common difference 12.

-2+9+20+31+42 = 100, is an arithmetic series with 1st term -2
 and common difference 11.

0+10+20+30+40 = 100, is an arithmetic series with 1st term 0
 and common difference 10.

2+11+20+29+38 = 100, is an arithmetic series with 1st term 2
 and common difference 9.

4+12+20+28+36 = 100, is an arithmetic series with 1st term 4
 and common difference 8.

6+13+20+27+34 = 100, is an arithmetic series with 1st term 6
 and common difference 7.

8+14+20+26+32 = 100, is an arithmetic series with 1st term 8
 and common difference 6.

10+15+20+25+30 = 100, is an arithmetic series with 1st term 10
 and common difference 5.

12+16+20+24+28 = 100, is an arithmetic series with 1st term 12
 and common difference 4.

14+17+20+23+26 = 100, is an arithmetic series with 1st term 14
 and common difference 3.

16+18+20+22+24 = 100, is an arithmetic series with 1st term 16
 and common difference 2.

18+19+20+21+22 = 100, is an arithmetic series with 1st term 18
 and common difference 1.

20+20+20+20+20 = 100, is an arithmetic series with 1st term 20
 and common difference 0.

22+21+20+19+18 = 100, is an arithmetic series with 1st term 22
 and common difference -1.

24+22+20+18+16 = 100, is an arithmetic series with 1st term 24
 and common difference -2.

26+23+20+17+14 = 100, is an arithmetic series with 1st term 26
 and common difference -3.

28+24+20+16+12 = 100, is an arithmetic series with 1st term 28
 and common difference -4.

30+25+20+15+10 = 100, is an arithmetic series with 1st term 30
 and common difference -5.

Edwin

Answer by Theo(13342)

i'll use an arithmetic series because there are formulas that should allow me to figure out what numbers could be used.




Sn = 1/2 * n * (A1 + An)

n = 5, therefore:

S5 = 1/2 * 5 * (A1 + A5)

multiply both sides by 2 to get:

2 * S5 = 5 * (A1 + A5)

divide everything by 5 to get:

2 * S5 / 5 = A1 + A5

subtract A1 from both sides of the equation to get:

2 * S5 / 5 - A1 = A5

let S5 equal 100 to get:

2 * 100 / 5 - A1 = A5

simplify to get:

40 - A1 = A5

since An = A1 + (n-1)*d, then:

40 - A1 = A1 + (n-1)*d

subtract A1 from both sides of this equation to get:

400 - 2A1 = (n-1)*d

distribute the multiplication to get:

40 - 2A1 = (n-1)*d

since n = 5, then:

40 - 2A1 = 4*d

solve for A1 to get:

A1 = 40 - 4*d)/2

simplify to get:

A1 = 20 - 2d

if this formula is good, then all we need to do is find a value for d that will allow us to subtract from 20 and get an integer (positive or negative should work). that value will be A1.

let's see if this works.

if d = 1, then A1 = 20 - 2 = 18.
A5 = A1 + 4d = 18 + 4 * 1 = 22
S5 = 1/2 * 5 * (18 + 22) = 1/2 * 5 * 40 = 1/2 * 200 = 100
it works !!!!!

if d = 2, then A1 = 20 - 4 = 16
A5 = 16 + 4*2 = 16 + 8 = 24
S5 = 1/2 * 5 * (16 + 24) = 1/2 * 5 * 40 = 1/2 * 200 = 100

id d = 4, then A1 = 20 - 8 = 12
A5 = 12 + 4*4 = 12 + 16 = 28
S5 = 1/2 * 5 * (12 + 28) = 1/2 * 5 * 40 = 1/2 * 200 = 100

it worked for all three !!!!

i used 3 different arithmetic series.
it was a matter of applying the formulas that come with arithmetic series.
An = A1 + (n-1)*d
Sn = n/2 * (A1 + An)

just to see if it will work with a negative number:

start with A1 = 20 - 2d.
let d = 100.
2d = 200
20 - 200 = -180
A1 = -180
d = 100
A5 = -180 + 4*100 = -180 + 400 = 220.
S5 = 5/2 * (-180 + 220) = 5/2 * 5*40/2 = 5*20 = 100

it works again !!!!!

the sequence is -180, -80, 20, 120, 220
add them up and you get 100.

you can probably do the same this with geometric series, and if you made up your own series rules, you might even be able to develop a formula using that.